Forrester and Cohen is a small accounting​ firm, managed by Joseph Cohen since the retirement in December of his partner Brad Forrester. Cohen and his 33 CPAs together bill 620 hours per month. When Cohen or another accountant bills more than 155 hours per​ month, he or she gets an additional​ "overtime" pay of​ $63.20 for each of the extra​ hours: this is above and beyond the 4,900 salary each draws during the month.​ (Cohen draws the same base pay as his​ employees.) Cohen strongly discourages any CPA from working​ (billing) more than 250 hours in any given month. The demand for billable hours for the firm over the next 6 months is estimated​ below:

Month

Estimate of Billable Hours

Jan.

610

Feb.

490

Mar.

1,020

Apr.

1,230

May

660

June

590

Cohen has an agreement with​ Forrester, his former​ partner, to help out during the busy tax​ season, up to 250 hours in any given month if​ needed, for an hourly fee of $125 Cohen will not even consider laying off one of his colleagues in the case of a slow economy. He​ could, however, hire another CPA at the same​ salary, as business dictates.

​a) Develop an aggregate plan for the​ 6-month period ​(enter your responses as whole​ numbers). Use regular​ time, then​ overtime, then​ Forrester, and then hire additional CPAs if needed.

​Note: For the CPA​ column, only include​ Cohen, his 3 ​CPAs, and any new CPAs he may hire in your total. Do NOT include Forrester.

Month

Estimate of Billable Hours

CPAs

Reg. time billable hours

Reg. time cost

​"Overtime" hours

​"Overtime" cost

Forrester hours

Forrester cost

Jan.

610

4

6206

​$19600

0

​$0

0

​$0

Feb.

490

Compute the cost of​ Cohen's plan of using overtime and Forrester.

The cost of​ Cohen's plan is ? (enter your response as a whole​ number).

​c) Should the firm remain as​ is, with a total of _____ CPAs?

A.

The firm should not remain as it is.

B.

The firm should remain as it is.

C.

One would have to carefully examine the other 6 months to see if hiring is merited
Answer #1

a) Aggregate plan based on strategy: regular time first, then overtime, then Forrester and then additional CPAs is following:

Reg. Time Billable Hours 620 620 620 620 620 620 Total 1 MonthEstimate of Reg.Time Overtime Overtime Forrester Forrester Cost

EXCEL FORMULAS:

Cell Formula Copy to
D2 =C2*155 D2:D7
E2 =C2*4900 E2:E7
F2 =MIN(C2*(250-155),MAX(0,B2-D2)) F2:F7
G2 =F2*63.2 G2:G7
H2 =MIN(250,MAX(0,B2-D2-F2)) H2:H7
I2 =H2*125 I2:I7
I11 =SUM(E2:E7,G2:G7,I2:I7)

b)

The cost of Cohen's plan is = $ 199,410

c) A. The firm should remain as it is, with a total of __4__ CPAs. (including Cohen)